# A new regularization scheme for the wave function of the Universe in the Lorentzian path integral

### v1

@@ I Introduction
The conformal problem can seemingly be avoided in the Lorentzian path integral formulation, since the integrand remains oscillatory even beyond the semiclassical level [10, 14]. Although the Lorentzian integral is only conditionally convergent, recent progress employing Picard–Lefschetz theory has made it possible to deform the integration contour onto steepest-descent thimbles, rendering the integral absolutely convergent and selecting the relevant semiclassical saddle points unambiguously [14] (see also Refs. [16, 17]). This approach yields a semiclassical exponent consistent with the tunneling wave function and leads to unsuppressed (inverse-Gaussian) fluctuations for linear perturbations [18] (see also Ref. [11] for an earlier study). Attempts to reproduce the HH wave function by introducing intrinsically complex lapse contours [19, 20, 21] inevitably lead to additional nonperturbative contributions, which again result in unsuppressed fluctuations [22, 23]. Alternative approaches impose suitable (Robin-type) boundary conditions on the perturbations [24, 25, 26] or on the scale factor [27, 28], aiming to recover either the tunneling or HH wave functions while maintaining well-behaved perturbations. Other proposals include the consideration of off-shell instantons in loop quantum cosmology [29, 30], analytic continuation from negative to positive potentials [31], and modifications of the dispersion relation motivated by trans-Planckian physics [32]. An alternative definition of the wave function based on saddle points satisfying regularity conditions has also been discussed in Ref. [33].
@@ II.1 Path integral for the scale factor
This boundary condition is of the Robin type, analogous to that introduced in Ref. [27]. Imposing $q_{\rm c}(1)=q_{1}$, the solution is
@@
$\int\frac{{\rm d}q_{0}}{\sqrt{2\pi}\,\sigma}\int_{q_{0}}^{q_{1}}\!\!\mathcal{D}q\;e^{\,i\,S_{g}^{\prime}}=\sqrt{\frac{c}{\pi i}}e^{\,i\,S_{g}^{\prime}[q_{c},N;q_{c}(0)]}\,.$ (16)
@@ III Path integral of the lapse function
$\tilde{\Psi}(q_{1},\sigma)\equiv\int_{0}^{\infty}{\rm d}N\sqrt{\frac{3\pi i}{2N}}e^{iS_{g}^{\prime}[N;q_{1},\sigma]}\,$, (30)
@@ IV.1 Saddle points
To obtain the HH wave function, we choose an original contour running from -$\infty$ to $\infty$ that passes below the singularity near the origin (located at $N\simeq-\,i\,3\pi^{2}\sigma^{2}$ for $\sigma>0$). The steepest-descent deformation in the lower half-plane then selects the two saddles with negative imaginary parts, $N_{\star}^{(\pm,-)}$ (see the red dotted curves in Fig. 3). In the limit $\sigma\to 0$, their combined contribution yields the HH wavefunction

### v2

@@ I Introduction
The conformal problem can seemingly be avoided in the Lorentzian path integral formulation, since the integrand remains oscillatory even beyond the semiclassical level [10, 14]. Although the Lorentzian integral is only conditionally convergent, recent progress employing Picard–Lefschetz theory has made it possible to deform the integration contour onto steepest-descent thimbles, rendering the integral absolutely convergent and selecting the relevant semiclassical saddle points unambiguously [14] (see also Refs. [16, 17]). This approach yields a semiclassical exponent consistent with the tunneling wave function and leads to unsuppressed (inverse-Gaussian) fluctuations for linear perturbations [18] (see also Ref. [11] for an earlier study). Attempts to reproduce the HH wave function by introducing intrinsically complex lapse contours [19, 20, 21] inevitably lead to additional nonperturbative contributions, which again result in unsuppressed fluctuations [22, 23]. Alternative approaches impose suitable (Robin-type) boundary conditions on the perturbations [24, 25, 26] or on the scale factor [27, 28, 29, 30, 31], aiming to recover either the tunneling or HH wave functions while maintaining well-behaved perturbations. Other proposals include the consideration of off-shell instantons in loop quantum cosmology [32, 33], analytic continuation from negative to positive potentials [34], and modifications of the dispersion relation motivated by trans-Planckian physics [35]. An alternative definition of the wave function based on saddle points satisfying regularity conditions has also been discussed in Ref. [36].
@@ II.1 Path integral for the scale factor
This boundary condition is of the Robin type, analogous to that introduced in Refs. [27, 28, 29, 30, 31]. Imposing $q_{\rm c}(1)=q_{1}$, the solution is
@@
Alternatively, it is instructive to evaluate the path integral in a more direct manner. First, we compute the classical contribution of the path integral $\int_{q_{0}}^{q_{1}}\mathcal{D}qe^{iS_{g}[q,N]}$. The classical solution $\tilde{q}_{c}$ with boundary conditions $\tilde{q}_{c}(0)=q_{0}$ and $\tilde{q}_{c}(1)=q_{1}$ is given by
$\tilde{q}_{c}(t)=H^{2}N^{2}t^{2}+\left(q_{1}-q_{0}-H^{2}N^{2}\right)t+q_{0}\,.$ (14)
Substituting this into the action, the corresponding classical action reads
$S_{g}[\tilde{q}_{c},N]$
$=-\frac{\pi^{2}}{2N}\left[3\left(q_{1}-q_{0}\right)^{2}+6N^{2}\left(H^{2}(q_{1}+q_{0})-2\right)+H^{4}N^{2}\right]\,.$ (15)
We then evaluate the integral over $q_{0}$ with the Gaussian weight $e^{-q_{0}^{2}/(2\sigma^{2})}$. Since the exponent is quadratic in $q_{0}$, the integral can be performed exactly:
$\int\frac{{\rm d}q_{0}}{\sqrt{2\pi}\,\sigma}e^{iS_{g}^{\prime}[\tilde{q}_{c},N;q_{0}]}=e^{iS_{g}^{\prime}}\sqrt{\frac{N}{N+i3\pi^{2}\sigma^{2}}}\,.$ (16)
The resulting classical action $S_{g}^{\prime}$ coincides with (LABEL:SgN). Moreover, substituting the center of the Gaussian, $q_{0}=i3\pi^{2}\sigma^{2}(q_{1}-H^{2}N^{2})/(N+i3\pi^{2}\sigma^{2})$, into (14) reproduces (12), which satisfies the boundary condition (11). We emphasize that the Robin boundary condition is not imposed from the beginning. Rather, the classical trajectory that dominates the path integral is automatically the one that satisfies this Robin boundary condition.
@@
We thus obtain
$\int\frac{{\rm d}q_{0}}{\sqrt{2\pi}\,\sigma}\int_{q_{0}}^{q_{1}}\!\!\mathcal{D}q\;e^{\,i\,S_{g}^{\prime}}=\sqrt{\frac{3\pi i}{2(N+i3\pi^{2}\sigma^{2})}}e^{\,i\,S_{g}^{\prime}[q_{c},N;q_{c}(0)]}\,.$ (19)
@@ III Path integral of the lapse function
$\tilde{\Psi}(q_{1},\sigma)\equiv\int_{0}^{\infty}{\rm d}N\sqrt{\frac{3\pi i}{2(N+i3\pi^{2}\sigma^{2})}}e^{iS_{g}^{\prime}[N;q_{1},\sigma]}\,$, (33)
@@ IV.1 Saddle points
To obtain the HH wave function, we choose an original contour running from -$\infty$ to $\infty$ that passes below the singularity near the origin (located at $N\simeq-\,i\,3\pi^{2}\sigma^{2}$ for $\sigma>0$). [^3] See Ref. [37] for an equivalent choice of contour discussed in a different context. The steepest-descent deformation in the lower half-plane then selects the two saddles with negative imaginary parts, $N_{\star}^{(\pm,-)}$ (see the red dotted curves in Fig. 3). In the limit $\sigma\to 0$, their combined contribution yields the HH wavefunction